# TriForce: Lossless Acceleration of Long Sequence Generation with Hierarchical Speculative Decoding

### v2

@@ 4.2 Hierarchical Speculation
1:Input: Prefix $[x_{1},\cdots,x_{t}]$, target model $M_{p}$ with full cache $C_{p}$, draft model $M_{q}$ with StreamingLLM cache $C_{q}$, target sequence length $T$, speculation length $\gamma_{1},\gamma_{2}$, auto-regressive drafting phase Draft, verification phase Verify, and correction phase Correct;
2:Initialize: Prefill $M_{p}$, $M_{q}$ with Prefix, construct retrieval cache $C_{r}$ using $x_{t}$, $N\leftarrow t$
@@
4.3 Algorithm
TriForce is devised to exploit the bottlenecks associated with both model weights and KV cache to enhance the inference speed of LLMs for long sequence generation. We present the pseudocode for the TriForce in Algorithm 1. It starts by prefilling the target model $M_{p}$ with full cache $C_{p}$ and draft model $M_{q}$ with StreamingLLM cache $C_{q}$ using a given input prefix, and then constructs the retrieval cache $C_{r}$. The initialization and update mechanism for the retrieval cache $C_{r}$ is guided by the insights of contextual locality discussed in Section 3.2. We first construct $C_{r}$ using the last token of the prefix, arranging tokens by their descending order of importance. In subsequent inferences, we overwrite the tokens with the least importance, maintaining the relevance and utility of the cache. A reconstruction of $C_{r}$ is triggered either when the rolling average acceptance rate drops below a threshold or at a designed stride.
The inference progresses iteratively until it reaches the target sequence length $T$. After each iteration, cache $C_{r}$ and $C_{q}$ are updated to prepare for the subsequent speculation phase. Each iteration encompasses two speculations: initially, $M_{q}$ utilizes $C_{q}$ to predict $M_{p}$ with $C_{r}$ for $\gamma_{1}$ steps until $n\geq\gamma_{2}$. Subsequently, these $n$ tokens are self-verified [51] by $M_{p}$ with $C_{p}$. This process constructs a hierarchy: the first layer of hierarchy employs a smaller, faster model $M_{q}$ with local context $C_{q}$ to speculate the large model $M_{p}$ with partial but high-quality global context $C_{r}$, addressing the model weights bottleneck. The second layer utilizes model $M_{p}$ with retrieval cache for self-speculation, overcoming the bottleneck caused by KV cache. This hierarchical speculation algorithm boosts efficiency by effectively addressing both bottlenecks.
System Implementation.
We implement the draft and target models using Transformers [44]. Leveraging a predetermined cache budget enables the effective use of PyTorch CUDA graphs [31, 30], significantly minimizing the kernel launching overhead during speculative decoding. FlashAttention is used to accelerate attention operations [11, 10, 21, 14]. Notably, we maintain full layer sparsity without omitting the initial two layers for system efficiency, ensuring our approach stays within the fixed KV cache budget. Although this strategy might lead to a lower acceptance rate, the benefit of maintaining a constant drafting cost makes it a worthwhile compromise. Additionally, to facilitate a faster speculation phase, we also implement two extra speculation cache structures, including our retrieval cache and StreamingLLM cache.
@@ 5 Empirical Evaluation
In this section, our goal is to showcase the capabilities of TriForce, a scalable and robust speculative decoding algorithm designed to expedite the inference of LLMs for long sequence generation, which significantly reduces the wall-clock time. We first present our end-to-end system, highlighting the overall speedup achieved, including both on-chip and offloading settings, followed by detailed ablation experiments.
  • In Section 5.1, we illustrate TriForce’s remarkable end-to-end performance, achieving up to 2.31$\times$ speedup for Llama2-7B-128K on an A100 and 7.78$\times$ on two RTX 4090 GPUs with offloading (achieving the latency as low as 0.108 s/token). We also examine the scalability and robustness of TriForce. Moreover, our results show that TriForce enhances large batch inference efficiency by 1.9$\times$.
  • In Section 5.2, we present detailed studies on the KV cache budget and chunk size, which provide valuable insights into selecting the appropriate hyperparameters to maximize efficiency. Additionally, we discuss the compatibility of TriForce with tree-based speculative decoding.
@@ 5.1 End-to-end Results
We now demonstrate that TriForce accelerates long sequence generation, up to 2.31$\times$ on an A100 GPU in the on-chip setting and 7.78$\times$ on two RTX 4090 GPUs with offloading for Llama2-7B-128K.
@@
Our experiments are based on Llama2 and LWM models with 128K context window size [43, 24, 32], which serve as our target models. In this setup, we utilize a 4K retrieval cache as an intermediate draft cache in our hierarchical system, while leveraging the JackFram/Llama68M (JF68M) [29] model as the initial draft model. For experiments involving offloading, we aim to maximize memory utilization by filling it up as much as possible and offloading the remaining KV cache to the CPU (AMD EPYC 9754 @ 2.25 GHz), while keeping the model weights on the GPU. Our evaluation is carried out on the PG-19 [34] and NarrativeQA [20] dataset, each testing on 100 examples, configured to a prompt length of 122K for on-chip settings and 127K for offloading settings, and aiming for a generation of 256 tokens. The performance of TriForce is analyzed across various hardware configurations, including on-chip experiments on an A100, and offloading experiments on RTX 4090 GPUs.
@@
Table 3: Offloading results (RTX 4090 GPUs): We present latency comparison between TriForce and Auto-regressive (AR) baseline for various models on different GPU setups. The sampling temperature is set to 0.6. The results indicate that TriForce achieves significant speedups across a range of models and hardware configurations. The entries marked with an asterisk represent the baseline using DeepSpeed-ZeRO-Inference [2].
@@
Figure 5: TriForce’s excellent scalability with longer contexts
@@
We evaluate TriForce using different temperatures, as depicted in Table 2. We observe that TriForce reaches up to 2.31$\times$ speedup for the on-chip setting with a minimal 4K KV cache budget for Llama2-7B-128K. For offloading settings, we provide end-to-end results on consumer GPUs for more models, including Llama2-7B-128K, Llama2-13B-128K, and LWM-Text-Chat-128K. Remarkably, in Table 3 we demonstrate that TriForce can efficiently serve a Llama2-13B with 128K contexts on two RTX 4090 GPUs, reaching an average time between tokens as low as 0.226 seconds, which is 7.94$\times$ faster than a highly optimized offloading system. Furthermore, with TriForce, Llama2-7B-128K can be served with 0.108s/token—only half as slow as the auto-regressive baseline on an A100. We also illustrate how TriForce boosts the efficiency of batched inference, a more frequently employed setting in real-world model serving. TriForce achieves 1.9$\times$ for a batch size of six, with each sample in the batch having 19K contexts, which is demonstrated in Table 4.
@@ 5.2.1 KV Cache Budget
As illustrated in Figure 6(a), for Llama2-7B-128K, there is a notable rise in the acceptance rate as the cache budget is increased to 4K, beyond which the rate starts to plateau, moving towards 1.0 gradually. This pattern indicates that while increasing the cache size up to a certain point can improve the acceptance rate, any further expansion beyond 4K may lead to diminishing benefits due to the increased latency of drafting. Consequently, setting the KV cache budget at 4K emerges as the most effective strategy for TriForce, ensuring a balance between achieving a high acceptance rate and minimizing additional drafting overhead.
Figure 6: (a) Analyzing speedup and acceptance rates across varying KV cache budgets reveals that a 4K budget is the optimal selection, balancing between acceptance rates and the overhead of drafting. (b) For 4K KV cache budget, it indicates that excessively small chunk sizes might overfit individual tokens, whereas overly large chunk sizes could constrain our diversity in selection. (c) TriForce is compatible with tree-based speculations such as Sequoia and Independent Sequences, enhancing the theoretical average number of tokens generated per decoding step of the target model by employing larger speculation budgets.
@@ 5.2.2 KV Cache Chunk Size
Since we utilized contextual locality to reuse the constructed retrieval cache, we need to examine how the choice of KV cache chunk size affects performance. Figure 6(b) shows that smaller chunk sizes may become too specialized, overfitting on a single token at the expense of generalization for future tokens, especially given the high attention sparsity in our methodology. Meanwhile, selecting chunk sizes that are too large can lead to the risk of high-score tokens being neutralized by surrounding low-score tokens, potentially resulting in a lack of differentiation among chunks. Such a strategy also limits our selection flexibility, constraining the diversity achievable within a fixed KV cache budget. For example, we are only allowed to select two chunks when the chunk size is set to 2K with 4K budget.
@@ 5.2.3 Compatibility with Tree-based Speculative Decoding
Tree-based speculative decoding methods [29, 41, 7, 5] leverage multiple candidates to enhance the acceptance rate so that more tokens can be accepted. We explore the possibility of integrating TriForce with tree-based speculative decoding. Specifically, for Llama2-7B-128K on an A100, we estimate the theoretical number of generated tokens when TriForce is combined with tree structures, including Sequoia and Independent Sequences [7]. As depicted in Figure 6(c), combining TriForce with tree-based speculation can potentially improve the overall end-to-end speedup by utilizing additional speculation budgets.
@@ 6 Conclusion
In this work, we introduced TriForce, a hierarchical speculative decoding system aimed at significantly enhancing the efficiency of serving LLMs with long contexts. Leveraging insights from attention sparsity and contextual locality, TriForce mitigates the dual bottlenecks associated with KV cache and model weights. Our empirical experiments demonstrate TriForce’s remarkable performance, including a notable speedup of up to 2.31$\times$ on A100 GPUs and an extraordinary 7.78$\times$ in offloading scenarios on two RTX 4090 GPUs, achieving 0.108s/token—only half as slow as the auto-regressive baseline on an A100. On a single RTX 4090 GPU, TriForce is 4.86$\times$ faster than DeepSpeed-Zero-Inference [2]. Additionally, it attains a 1.9$\times$ speedup with large batches. These achievements illustrate TriForce’s potential to revolutionize the serving of long-context models for long sequence generation.

### v3

@@ 4.3 Algorithm
1:Input: Prefix $[x_{1},\cdots,x_{t}]$, target model $M_{p}$ with full cache $C_{p}$, draft model $M_{q}$ with StreamingLLM cache $C_{q}$, target sequence length $T$, speculation length $\gamma_{1},\gamma_{2}$, drafting phase Draft, verification phase Verify, and correction phase Correct;
2:Initialize: Prefill $M_{p}$, $M_{q}$, construct retrieval cache $C_{r}$ using $x_{t}$, $N\leftarrow t$
@@ 5 Empirical Evaluation
In this section, our goal is to showcase the capabilities of TriForce, a scalable and robust speculative decoding algorithm designed to expedite the inference of LLMs for long sequence generation, which significantly reduces the wall-clock time. We first present our end-to-end system, highlighting the overall speedup achieved, including both on-chip and offloading settings, followed by a comparison with other methods and ablation experiments.
@@ 5.1 End-to-end Results
We demonstrate that TriForce accelerates long sequence generation, up to 2.31$\times$ on an A100 in the on-chip setting and 7.78$\times$ on two RTX 4090s with offloading for Llama2-7B-128K.
@@
Our experiments are based on Llama2 and LWM models with 128K context window size (Touvron et al., 2023; Liu et al., 2024a; Peng et al., 2023), which serve as our target models. In this setup, we utilize a 4K retrieval cache as an intermediate draft cache in our hierarchical system, while leveraging the JackFram/Llama68M (JF68M) (Miao et al., 2023) model as the initial draft model. For experiments involving offloading, we aim to maximize memory utilization by filling it up as much as possible and offloading the remaining KV cache to the CPU (AMD EPYC 9754 @ 2.25 GHz), while keeping the model weights on the GPU. Our evaluation is carried out on the PG-19 (Rae et al., 2019) and NarrativeQA (Kočiský et al., 2018) dataset, each testing on 100 examples, configured to a prompt length of 122K for on-chip settings and 127K for offloading settings, and aiming for a generation of 256 tokens. The performance of TriForce is analyzed across various hardware configurations, including on-chip experiments on an A100, and offloading experiments on RTX 4090 GPUs.
@@
Table 3: Offloading results (RTX 4090): We present latency comparison between TriForce and Auto-regressive (AR) baseline for various models on different GPU setups. The sampling temperature is set to 0.6. The results indicate that TriForce achieves significant speedups across a range of models and hardware configurations. The entries marked with an asterisk represent the baseline using DeepSpeed-ZeRO-Inference (Aminabadi et al., 2022).
@@
Figure 5: TriForce’s excellent scalability with longer contexts
@@
We evaluate TriForce using different temperatures, as depicted in Table 2. We observe that TriForce reaches up to 2.31$\times$ speedup for the on-chip setting with a minimal 4K KV cache budget for Llama2-7B-128K. For offloading settings, we provide end-to-end results on consumer GPUs for more models, including Llama2-7B-128K, Llama2-13B-128K, and LWM-Text-Chat-128K. Remarkably, in Table 3 we demonstrate that TriForce can efficiently serve a Llama2-13B with 128K contexts on two RTX 4090s, reaching an average time between tokens as low as 0.226 seconds, which is 7.94$\times$ faster than a highly optimized offloading system. Moreover, with TriForce, Llama2-7B-128K can be served with 0.108s/token—only half as slow as the auto-regressive baseline on an A100. We also illustrate how TriForce boosts the efficiency of batched inference, a more frequently employed setting in real-world model serving. TriForce achieves 1.9$\times$ for a batch size of six, with each sample in the batch having 19K contexts, which is demonstrated in Table 4.
@@
Method Speedup TriForce 2.31$\times$ REST 1.47$\times$ Skipping Layers 1.36$\times$
Table 5: Speedup comparison with REST and Skipping Layers
5.2 Comparison with Other Methods
We provide a comparison with REST (He et al., 2023) and Skipping Layers (Zhang et al., 2023). Table 5 compares TriForce, REST, and Skipping Layers with Llama2-7B-128K on an A100 using PG-19 dataset, showing TriForce achieves the best speedup for long sequence generation.
TriForce retrieves information from the KV cache, allowing dynamic adaptation to contexts, while REST uses an external predefined datastore. In Table 5, Skipping Layers utilizes 68% of the KV cache, whereas TriForce efficiently uses only 3%, addressing the bottleneck in long-context scenarios better.
@@ 5.3.1 KV Cache Budget
As illustrated in Figure 6(a), for Llama2-7B-128K, the acceptance rate rises with the cache budget up to 4K, then plateaus towards 1.0. This suggests that increasing the cache size beyond 4K offers diminishing benefits due to drafting latency. Thus, a 4K KV cache budget is optimal for TriForce, balancing high acceptance rates and minimal drafting overhead.
Figure 6: (a) Analyzing speedup and acceptance rates across different KV cache budgets reveals that a 4K budget is optimal, balancing acceptance rates and the drafting overhead. (b) For a 4K KV cache budget, excessively small chunk sizes may overfit individual tokens, while overly large chunk sizes could limit selection diversity. (c) TriForce is compatible with tree-based speculations, enhancing the theoretical average number of tokens generated per decoding step of the target model by employing larger speculation budgets.
@@ 5.3.2 KV Cache Chunk Size
Since we utilize contextual locality to reuse the retrieval cache, we need to examine the impact of KV cache chunk size on performance. Figure 6(b) shows that smaller chunks may overfit to single tokens, limiting generalization, while larger chunks may dilute high-score tokens with low-score ones, resulting in reduced differentiation among chunks. Large chunks also reduce selection flexibility, constraining diversity within a fixed cache budget.
@@ 5.3.3 Compatibility with Tree-based Speculative Decoding
We explore the possibility of integrating TriForce with tree-based speculative decoding. Specifically, for Llama2-7B-128K on an A100, we estimate the theoretical number of generated tokens when TriForce is combined with tree structures, including Sequoia (Chen et al., 2024) and Independent Sequences. As depicted in Figure 6(c), this integration can potentially improve the end-to-end speedup by utilizing additional speculation budgets.
@@ 6 Conclusion
In this work, we introduced TriForce, a hierarchical speculative decoding system aimed at significantly enhancing the efficiency of serving LLMs with long contexts. Leveraging insights from attention sparsity and contextual locality, TriForce mitigates the dual bottlenecks associated with KV cache and model weights. Our empirical experiments demonstrate TriForce’s remarkable performance, including a notable speedup of up to 2.31$\times$ on an A100 and 7.78$\times$ on two RTX 4090s with offloading, achieving 0.108s/token—only half as slow as the auto-regressive baseline on an A100. These achievements illustrate TriForce’s potential to revolutionize the serving of long-context models for long sequence generation.
@@ Acknowledgments
Appendix A System Implementation
We implement the draft and target models using Transformers (Wolf et al., 2019). Leveraging a predetermined cache budget enables the effective use of PyTorch CUDA graphs (Paszke et al., 2019; NVIDIA & Fitzek, 2020), significantly minimizing the kernel launching overhead during speculative decoding. FlashAttention is used to accelerate attention operations (Dao et al., 2022; Dao, 2023; Kwon et al., 2023; Hong et al., 2023). Notably, we maintain full layer sparsity without omitting the initial two layers for system efficiency, ensuring our approach stays within the fixed KV cache budget. Although this strategy might lead to a lower acceptance rate, the benefit of maintaining a constant drafting cost makes it a worthwhile compromise. Additionally, to facilitate a faster speculation phase, we also implement two extra speculation cache structures, including our retrieval cache and StreamingLLM cache.
Appendix B Additional Experiments
B.1 TriForce’s Scalability for Longer Inputs
For longer inputs, a solution is offloading the KV cache to the CPU memory. Below are experiments on a single L40 GPU with offloading for longer inputs on LWM-Text models (Liu et al., 2024a) using the PG-19 dataset, showing impressive speedup.
Model Speedup
LWM-Text-256K 11.81$\times$
LWM-Text-512K 12.10$\times$
Table 6: TriForce’s Scalability for longer inputs
B.2 TriForce’s Scalability for Longer Outputs
In the previous sections of the paper, we focused on a long input and a short output to illustrate the efficiency gains in a commonly encountered setting. Here we provide additional experiments with longer output sequences up to 2K with Llama2-7B-128K on an A100 GPU using the PG-19 dataset:
Output Sequence Length Speedup
256 2.31$\times$
512 2.28$\times$
768 2.34$\times$
1024 2.32$\times$
1536 2.28$\times$
2048 2.29$\times$
Table 7: TriForce’s Scalability for longer outputs
As seen in Table 7, the performance of TriForce remains relatively stable. This stability is due to the retrieval cache being reconstructed at fixed intervals, ensuring it stays aligned with the evolving context. Furthermore, our retrieved cache can be effectively reused across multiple decoding steps, amortizing the overhead of retrieval. These results demonstrate the scalability of TriForce and its robustness in handling varying output lengths, thereby addressing the concern about its generalizability.
B.3 Ablation of $\gamma_{1},\gamma_{2}$
Here we provide additional ablation results of $\gamma_{1},\gamma_{2}$ in Table 8 with Llama2-7B-128K on an A100 GPU using the PG-19 dataset. Due to the gap between JF68M with StreamingLLM and Llama-7B-128K with retrieval cache, the acceptance rate of the first speculation phase in our speculation hierarchy is low. As shown in the table, the optimal value for $\gamma_{1}$ is 2 and for $\gamma_{2}$ is 6.
$\gamma_{1}$ Speedup
1 2.30$\times$
2 2.31$\times$
3 2.24$\times$
4 2.17$\times$
5 2.09$\times$
6 2.01$\times$
 7 1.92$\times$
8 1.83$\times$
(a) Speedup at different $\gamma_{1}$ values with $\gamma_{2}=6$
$\gamma_{2}$ Speedup
1 1.62$\times$
2 1.90$\times$
3 2.13$\times$
4 2.22$\times$
5 2.26$\times$
6 2.31$\times$
7 2.24$\times$
8 2.20$\times$
(b) Speedup at different $\gamma_{2}$ values with $\gamma_{2}=2$
Table 8: Ablation of $\gamma_{1},\gamma_{2}$